# Intent3D: 3D Object Detection in RGB-D Scans Based on Human Intention

### v2

@@ 5.3 Evaluation Metrics
We use Top1-Accuracy and Average Precision for evaluation. Top1-Accuracy gauges the accuracy of the model’s highest confidence prediction, while Average Precision evaluates precision across varying confidence thresholds. We calculate the Intersection over Union (IoU) between predicted and ground truth boxes with thresholds of 0.25 and 0.5 to indicate correctness.
5.4 Quantitative Comparisons
Results of our IntentNet Due to explicit modeling the comprehension of intention language and reasoning on box candidates with cascaded optimization, our IntentNet outperforms all the previous methods by a large margin, achieving 11.22%, 8.05% improvement on Top1-Acc@0.25 and Top1-Acc@0.5 and 9.12%, 5.43% improvement on AP@0.25 and AP@0.5 on the val set, compared with the second best method. Similarly, on the test set, our IntentNet outperforms all the previous methods by a large margin, achieving 11.06%, 10.84% improvement on Top1-Acc@0.25 and Top1-Acc@0.5 and 6.72%, 5.6% improvement on AP@0.25 and AP@0.5 compared with the second best method.
Results of expert models The expert models mostly focus on aligning with nouns on the sentence, since they are originally built for targeting referential language. However, noun is not the target but just one of the object for a certain verb in our intention language. Therefore, they perform inferiorly compared with our IntentNet, where we explicitly comprehend all the verb-object parts. It worth notice that a better 3D-VG model, EDA, performs worse than its baseline, BUTD-DETR, since EDA further digs into the alignment with the attributes or auxiliary object toward the noun in the sentence, which is not the target in our case and thus misleads the model.
Results of foundation model The foundation model, 3D-VisTA, performs better than the expert models on most metrics, which is attributed to its widely multimodal alignment from pretraining. However, it still falls short of our IntentNet, since their full reliance on the detection outputs from the detector [49], which are not perfect on evaluation. Instead, we only take the predictions from the pretrained detector as visual reference and provide reasoning on those box candidates. Therefore, even if we are using a less powerful detector [38] than theirs, we still outperform them.
Results of LLM-based model The LLM-based model, Chat-3D-v2, performs the worst in both zero-shot and finetuning on most metrics, which is expected given that current LLM-based models perform poorly on 3D-VG [16] and 3D-IG is even more challenging. The main bottleneck might still be the lack of training data for aligning 3D visual features into LLM space [16, 18]. Also, due to hallucination problems in the LLM, Chat-3D-v2 performs significantly worse on AP metrics, which are more sensitive to the false positive predictions. Its decent performance on Top1-Acc is mainly due to the use of a pre-trained detector [21] and the powerful LLM to reason out the correct category of the target.
Table 3: Ablation studies on Intent3D’s val set. "Verb" indicates the alignment with verb tokens in the sentence, which is learned by $L_{vPos}$ and $L_{vSem}$. "Verb2Obj" indicates the alignment with object token in the sentence when the queries are modulated by the verb. "MatchBox" indicates the Candidate Box Matching, "Adapt" indicates the Cascaded Adaptive Learning.
@@ 5.5 Ablation Studies
Ablation studies are conducted on the val set to evaluate the effectiveness of each component in our IntentNet. As shown in Tab. 3, we apply controlling variables, removing one component in each experiment based on the comparison with the complete verison (e). We observe that every component is essential for optimal performance, as the exclusion of any component results in a notable decline in performance. Specifically, the comparison between (a) and (e) shows that the alignment with verb tokens in the sentence is foundamental for the model to understand the intention, which is learned by $L_{vPos}$ and $L_{vSem}$. Omitting this causes a significant reduction in performance. The contrast between (b) and (e) illustrates that additionally learning about reasoning about the object given verb can further improve the comprehension of the intention. The evaluation between (c) and (e) demonstrates that the Candidate Box Matching effectively extracts distinctive features near the target boxes to boost overall performance. Lastly, the comparison between (d) and (e) confirms that Cascaded Adaptive Learning facilitates the model to concurrently optimize each of the objectives.
Figure 5: Qualitative results of ablation studies. "Verb" indicates the alignment with verb tokens. "Verb2Obj" indicates the alignment with object token in the sentence when the queries are modulated by the verb. "MatchBox" indicates the Candidate Box Matching, "Adapt" indicates the Cascaded Adaptive Learning. Red boxes indicate the ground truth, Blue boxes indicate the counterpart model’s prediction, and Green boxes indicate our model’s prediction.
@@ 5.6 Qualitative Results
We provide qualitative results of ablation studies in Fig. 5 to further demonstrate the effectiveness of each component in our IntentNet. The first column: The model without alignment with the verb fails to grasp the intention to identify the target. This highlights the fundamental importance of aligning with verb tokens. The second column: The model lacking alignment with object tokens when given verb tokens does not fully capture the intention. In this context, "add" is too vague to specifically identify "plant". However, with a better understanding of the object "greenery", our final model can accurately detect the intended object. The third column: The model without the Candidate Box Matching cannot effectively extract distinctive box features near the target, leading it to identify irrelevant items. The fourth column: The model lacking Cascaded Adaptive Learning does not efficiently optimize each objective. Hence, despite understanding the intention and reasoning about the target "chair" through contrastive learning, it fails to localize all instances of the target due to insufficient optimization of the regression loss.
@@ 6 Conclusion
In this paper, we propose a new challenging task, 3D Intention Grounding, designed to detect target objects based on human intention within 3D scenes. We introduce a new dataset, Intent3D, with detailed statistics to support this task. To tackle this problem, we first construct comprehensive baselines involving current expert model, foundational model, and LLM-based model. Finally, we propose a novel method, IntentNet, which includes three key aspects: intention comprehension, reasoning to identify object candidates, and cascaded optimization of objective functions, to effectively perform 3D Intention Grounding.

### v3

@@ 5.3 Evaluation Metrics
We use Top1-Accuracy and Average Precision for evaluation. Top1-Accuracy measures the accuracy of the model’s highest-confidence prediction, while Average Precision evaluates precision across varying confidence thresholds. We compute the Intersection over Union (IoU) between predicted and ground truth boxes with thresholds of 0.25 and 0.5 to indicate correctness.
Table 3: Ablation studies on Intent3D’s val set. “Verb” indicates the alignment with verb tokens in the sentence, which is learned by $L_{vPos}$ and $L_{vSem}$. “Verb2Obj” indicates the alignment with object token in the sentence when the queries are modulated by the verb. “MatchBox” indicates the Candidate Box Matching, “Adapt” indicates the Cascaded Adaptive Learning.
@@
5.4 Quantitative Comparisons
Results of our IntentNet Due to the explicit modeling of intention language comprehension and reasoning on box candidates with cascaded optimization, our IntentNet significantly outperforms all previous methods. Compared to the second-best approach on the validation set, it achieves improvements of 11.22% and 8.05% in Top1-Acc@0.25 and Top1-Acc@0.5, respectively. Additionally, it improves AP@0.25 and AP@0.5 by 9.12% and 5.43%, respectively. Similarly, on the test set, we obtain 11.06%, 10.84% improvement in Top1-Acc@0.25 and Top1-Acc@0.5, respectively; and 6.72%, 5.6% improvement in AP@0.25 and AP@0.5, respectively.
Results of expert models The expert models mostly focus on aligning with nouns in the sentence, since they were originally built for targeting referential language. However, in our intention language, the noun is not the target but only one aspect of an object associated with a certain verb. Therefore, they perform inferiorly compared with our IntentNet, where we explicitly comprehend all the verb-object relations. Notably, a better 3D-VG model, EDA, performs worse than its baseline, BUTD-DETR, since EDA additionally focuses the alignment with the auxiliary object related to the noun in the sentence, which is not the target in our case and thus misleads the model.
Results of foundation model The foundation model, 3D-VisTA, performs better than the expert models on most metrics, which is attributed to its widely multimodal alignment in pretraining. However, it still falls short of our IntentNet because it fully relies on detection outputs from the detector [Schult et al., 2022], which are not perfect for evaluation. Instead, we only take the predictions from the pretrained detector as visual reference and provide reasoning on those box candidates. Therefore, even if we are using a less powerful detector [Liu et al., 2021b], we still outperform them.
Results of LLM-based model The LLM-based model, Chat-3D-v2, performs the worst in both zero-shot and finetuning on most metrics, which is expected given that current LLM-based models perform poorly on 3D-VG [Hong et al., 2024] and 3D-IG is even more challenging. The main bottleneck might still be the lack of training data for aligning 3D visual features into LLM space [Hong et al., 2024; Huang et al., 2023]. Also, due to hallucination problems in the LLM, Chat-3D-v2 performs significantly worse on AP metrics, which are more sensitive to the false positive predictions. Its decent performance on Top1-Acc is mainly due to the use of a pretrained detector [Jiang et al., 2020] and the powerful LLM to reason out the correct category of the target.
An additional comparison with GPT-4 is elaborated in Appendix A.6.
Figure 5: Qualitative results of ablation studies. “Verb” indicates the alignment with verb tokens. “Verb2Obj” indicates the alignment with object token in the sentence when the queries are modulated by the verb. “MatchBox” indicates the Candidate Box Matching, “Adapt” indicates the Cascaded Adaptive Learning. Red boxes indicate the ground truth, Blue boxes indicate the counterpart model’s prediction, and Green boxes indicate our model’s prediction.
@@ 5.5 Ablation Studies
Ablation studies are conducted on the val set to evaluate the effectiveness of each component in our IntentNet. As shown in Tab. 3, we apply controlled variables, removing one component at a time for comparison with the complete version (e). We observe that every component is essential for optimal performance, since the exclusion of any component results in a notable decline in performance. Specifically, the comparison between (a) and (e) shows that the alignment with verb tokens in the sentence is fundamental for the model to understand the intention, which is learned by $L_{vPos}$ and $L_{vSem}$. Omitting this causes a significant reduction in performance. The comparison between (b) and (e) shows that incorporating reasoning about the object, given the verb, further enhances the model’s understanding of the intention. The evaluation between (c) and (e) demonstrates that the Candidate Box Matching module effectively extracts distinctive features near the target boxes to boost overall performance. Lastly, the comparison between (d) and (e) confirms that Cascaded Adaptive Learning facilitates the model to concurrently optimize each of the objectives.
@@ 5.6 Qualitative Results
We provide qualitative results of ablation studies in Fig. 5 to further demonstrate the effectiveness of each component in our IntentNet. The first column: The model without alignment with the verb fails to grasp the intention to identify the target. This highlights the fundamental importance of aligning with verb tokens. The second column: The model lacking alignment with object tokens when given verb tokens does not fully capture the intention. In this context, “add” is too vague to specifically identify “plant”. However, with a better understanding of the object “greenery”, our final model can accurately detect the intended object. The third column: The model without the Candidate Box Matching fails to effectively extract distinctive box features near the target, leading it to identify irrelevant items. The fourth column: The model lacking Cascaded Adaptive Learning does not efficiently optimize each objective. Hence, despite understanding the intention and reasoning about the target “chair” through contrastive learning, it fails to localize all instances of the target due to insufficient optimization of the regression loss.
@@ 6 Conclusion
We propose a new task, 3D Intention Grounding (3D-IG), designed to detect target objects based on human intention within 3D scenes. We introduce a new dataset, Intent3D, with detailed statistics to support this task. To tackle this problem, we first construct comprehensive baselines involving current expert model, foundational model, and LLM-based model. Finally, we propose a novel method, IntentNet, which includes three key aspects: intention comprehension, reasoning to identify object candidates, and cascaded optimization of objective functions, to tackle 3D-IG.
Acknowledgments: This research is supported by NSF IIS-2309073, ECCS-2123521 and Cisco Research unrestricted gift. This article solely reflects the opinions and conclusions of its authors and not the funding agencies.
Appendix A Appendix
A.1 Additional Discussion of Intent3D Data
Failed Cases Here, we aim to provide more details of failed cases in the responses from GPT. As shown in Table 4, (1) GPT sometimes fails to generate a complete sentence. (2) GPT may generate a sentence that includes Unicode characters. (3) GPT may generate a sentence that is not directly related to the object, since the name of the object’s category is polysemous. For instance, “keyboard” can refer to a musical instrument, but in our dataset, it actually denotes a computer peripheral. (4) We exclude objects whose categories are too general, such as “machine”, in which case GPT fails to generate a specific intentional text.
Table 4: Examples of failed cases in the responses from GPT.
(1) “bathroom cabinet” : I
[…]
(3) “keyboard” : I need to compose music for my upcoming online performance
(4) “machine” : I want to utilize the object (depending on the type) to assist me in tasks
API Usage In employing ChatGPT (GPT-4), adjustment of the temperature parameter enables control over output randomness. Higher temperature values foster greater diversity and creativity, while lower values yield more deterministic but potentially repetitive responses. We set temperature=1.2 to optimize the diversity of model outputs. We have explored alternative GPT versions such as gpt-3.5-turbo-1106, and gpt-4-1106-preview. However, empirical evaluation revealed that GPT-4 consistently outperforms others, making it our preferred choice.
Limitations The quality and variety of intention sentences generated for each object class depend on the performance of GPT-4. Although GPT-4 demonstrates strong proficiency in real-world knowledge, and we manually refined outputs to address garbled cases, some inherent subjectivity may still be introduced due to GPT-4’s training data.
A.2 Additional Implementation Details
Here, we provide additional details on the implementation of each method. For BUTD-DETR [Jain et al., 2022], we adhere to its official configuration in ScanRefer [Chen et al., 2020]. The batch size is set to 24, and the learning rate (which is the same as ours) decreases to one-tenth at the 65th epoch. We follow the suggestion in its source code to continue training the model while monitoring the validation accuracy until the model converges. It finally takes 100 epochs to converge. For EDA [Wu et al., 2023], we also follow its official configuration, where the batch size is set to 48. The learning rate for backbones is 0.002, and for the rest, it is 0.0002. The learning rate decreases to one-tenth at the 50th and 75th epochs. It takes 104 epochs to converge. For 3D-VisTA [Zhu et al., 2023], we use its official configuration in ScanRefer [Chen et al., 2020], where the batch size is set to 64, and the learning rate is 0.0001. The warm-up step is set to 5000. Although the official training epoch is 100, we find that the model converges at the 47th epoch. We adopt its pre-trained checkpoint provided in the source code for fine-tuning. In the case of Chat-3D v2, we follow its official instructions to prepare the annotations. It takes 3 epochs to fine-tune the model.
Figure 6: Visualization of correct cases and false cases of IntentNet. Red boxes are ground truth boxes, green boxes are prediction outputs
A.3 Additional Qualitative Results of IntentNet
In this section, we present additional qualitative results of our IntentNet, showcasing both correct and false cases.
Correct Cases In the upper row (Fig. 6(a)), IntentNet demonstrates impressive accuracy in various scenarios where the number of target instances exceeds one and the instances are not closely situated, as observed in cases of “cabinet” or “refrigerator”. Furthermore, for the less common intention “enjoy aesthetical pleasure”, IntentNet also performs well.
False Cases In the lower row (Fig. 6(b)), we highlight cases where IntentNet fails. Particularly, in the case of “washing machines”, IntentNet struggles to detect all instances in large environments with multiple targets (four instances). However, we also find that some false cases are attributed to the ambiguity or subjectivity in the bounding box annotations provided by ScanNet [Dai et al., 2017]. For instance, in the “desk” case, the annotation includes a double-person desk, while IntentNet detects only one of the single-person desks, which is also reasonable.
A.4 Additional Discussion of Ambiguity
In our dataset, intentions and objects are not one-to-one association. For example, “I want to sleep” relates to both “couch” in living room and “bed” in bedroom. However, the reason we generate intention towards a “bed” in a “bedroom” that does not have a “couch” in it, is to provide a clear and unambiguous mapping of intention to object for better training. As shown in Fig. 7, we input “I intend to rest my legs” with the above ambiguous scenario where multiple objects can fulfill this intention. When we lower the confidence score filter to allow multiple outputs, our model can ground many reasonable objects (“footrest”, “couch”, “table”, etc.), since these objects have similar intentions in our training set but just appear in different scenes for unambiguous training.
Figure 7: Example for multiple objects prediction.
A.5 Additional Discussion of Human-likeness
To assess the concern regarding whether ChatGPT-generated sentences exhibit human-like linguistic naturalness, similar to [Qu et al., 2024], we randomly selected 500 original responses produced by ChatGPT and asked three PhD students to manually evaluate them. The results show that all intention sentences meet the criteria for human-like quality.
A.6 Additional Comparison with GPT-4
Table 5: Compare with two-stage method based on GPT-4.
Method Detector top1-acc@0.25 top1-acc@0.5 ap@0.25 ap@0.5
GPT-4 GroupFree 41.40 28.40 15.10 7.76
IntentNet GroupFree 58.34 40.83 41.90 25.36
We provide an additional comparison with a two-stage method based on GPT-4. For the two-stage method, we fairly use the same detector as ours to predict bounding boxes and categories. Given the intention and the detector’s predictions, GPT-4 selects the most relevant objects. Results are shown on validation set in Table 5. Due to the detector’s prediction errors and hallucinations from GPT-4, it performs poorly in AP metric.
A.7 Cases Where IntentNet Fails but Others fed with objects Succeed
We provide failed cases of our IntentNet to offer potential improvement directions for future work. We find that IntentNet sometimes struggles with intentions involving multiple actions.
Figure 8: A case where IntentNet fails (blue box) but EDA fed with ground truth object candidates succeeds (green box). The input intention is “I need to cool down on hot days while sitting in the living room.”
As shown in Fig. 8, consider the intention: “I need to cool down on hot days while sitting in the living room.” In this case, the primary intention is “cool down”, which targets the object “fan”. However, the secondary action “sitting” acts as a distractor, leading our model to incorrectly detect “couch” (blue bounding box). When we directly provide ground truth object to EDA, the task becomes simplified, allowing it to correctly detect the “fan” (green bounding box). Another example is shown in Fig. 9. The input intention is “I want to sit comfortably while doing my office work.” Here, “doing my office work” distracts IntentNet from correctly detecting the optimal object, causing it to wrongly identify “table” instead of “chair”, which should be the correct object for “sit comfortably”.
Figure 9: A case where IntentNet fails (blue box) but EDA fed with ground truth object candidates succeeds (green box). The input intention is “I want to sit comfortably while doing my office work.”
A.8 Visualization of grounding for trivial objects
In our data construction, we filter out trivial objects whose number of instances is greater than or equal to six, to increase the dataset’s difficulty and quality. However, we want to make it clear that this filtering is independently applied to each scene itself, not to the entire dataset. Specifically, as shown in Fig. 10, in this scene, “table” is a trivial object (with quantity greater than or equal to 6) and is filtered out during the training phase. But this does not mean that the entire dataset lacks the intention of “table”, because in other scenes, “table” may not be numerous, and therefore the intention regarding “table” still exists in the dataset. Thus, in the inference stage, the model is able to detect such common used objects (green boxes). Fig. 11 is another example, in which chairs are trivial objects.
Figure 10: A case where IntentNet grounds trivial objects (green box). The input intention is “I need to have a space where I can spread out and study.”
Figure 11: A case where IntentNet grounds trivial objects (green box). The input intention is “I want to sit down during long conferences.”
A.9 Detailed statistics
We provide detailed statistics of verb and noun toward each object category in Fig. 12 and Fig. 13, respectively. In Fig. 14, we also provide a detailed histogram of different objects we use in constructing our intention sentences. Here, every third data point is displayed for the sake of clarity.
Figure 12: Detailed statistics of verb used in intentions toward each object category
Figure 13: Detailed statistics of noun used in intentions toward each object category
Figure 14: The statistic of object type histogram.
A.10 Monitor the training process of IntentNet
We provide the accuracy curves and the AP curves of the training process of IntentNet. As shown in Fig. 15 and Fig. 16, the accuracy and AP of IntentNet both converge and achieve strong performances, which indicates that our model trains effectively by the training set.
Figure 15: Accuracy curves of the training process of IntentNet.
Figure 16: AP curves of the training process of IntentNet.
A.11 Details of IntentNet
As shown in Fig. 17, We provide a detailed figure to illustrate the connections in the network and key modules of our loss functions.
Figure 17: Details of IntentNet. K, V indicate the Key and Value of cross-attention layer, respectively.